Find the equation of the line.
Thru (4,2) ; m= -2/3
Thru (9,2) ; perpendicular to 3x+2y=31

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Line thru (4,2) where m= -2/3
y = (-2/3)x + b |using ordered pair Pt(4,2) to solve for b
2 = -8/3 + b
14/3 = b
y = (-2/3)x + 14/3 |standard form: 2x + 3y = 14
Thru (9,2)and 'perpendicular' to 3x+2y=31 OR y = (-3/2)x +31/2
slope of line perpendicular to it would be 2/3. (negative reciprocal)
y = (2/3)x + b |using ordered pair Pt(9,2) to solve for b
2 = 6 + b
-4 = b
y = (2/3)x -4